# HILCodec: High Fidelity and Lightweight Neural Audio Codec

### v1

@@ VI-C Computational Complexity
TABLE III: Complexity of various neural audio codecs.
@@
We compared HILCodec against other neural audio codecs in terms of computational complexity. We measured the number of parameters, multiply-add counts (MAC) for coding one second long input waveform, and real-time factor (RTF). RTF is defined as the ratio of the length of the input waveform to the time consumed for coding. An RTF higher than 1 means the codec operates in real-time. When calculating RTF, we used one thread of a server CPU (Intel Xeon 6248R 3GHz), exported each model to ONNX Runtime [69], and simulated a streaming condition. Namely, for HILCodec and EnCodec, a chunk with 320 waveform samples was encoded and decoded at a time, and for AudioDec, a chunk with 300 samples was processed at a time. For every convolution layer with a kernel size larger than one, previous samples were stored in a cache and used to calculate the current chunk. Descript and HiFi-Codec are not streamable, so their RTFs are not reported. The results are shown in Table III. HILCodec has the smallest number of parameters, requiring minimum memory to store and load model parameters. Moreover, the MAC of HILCodec is smaller than AudioDec, Descript, and HiFi-Codec, and its RTF is higher than 1.
@@ VII Conclusion
Appendix A Variance of Residual Branches
@@
In Section III-B2, we construct the VCRB such that the variance after initialization follows (6). However, in Fig. 2, although the variance increases linearly with the network depth, it does not strictly follow (6). It is demonstrated that for exactness, we must use scaled weight standardization (Scaled WS) [62]. However, in our early experiments, using Scaled WS resulted in a slightly worse ViSQOL score. Reference [62] also experienced similar degradation and hypothesized that Scaled WS harms the performance of depthwise convolutions. Therefore, as an alternative, we carefully apply initialization for each layer. For layers where an activation function does not come immediately after, we apply LeCun initialization so that $\mathbf{Var}(W\mathbf{x})\approx\mathbf{Var}(\mathbf{x})$ where $W$ is a weight parameter. For layers with subsequent activation functions, we apply He initialization so that $\mathbf{Var}(\sigma(W\mathbf{x})))\approx\mathbf{Var}(\mathbf{x})$. Through these methods, despite the model do not exactly follow (6), it achieves a linear variance increment.
Appendix B Differences Between MFBD and MSD
@@
We present detailed comparisons between SBD used in Avocodo and MFBD used in HILCodec. First, SBD decomposes the input signal into 16 sub-bands while MFBD uses co-prime numbers of sub-bands $N\in\{1,2,3,5,7,11\}$. As described in Section IV-A2, this enables MFBD to model both long- and short-term dependencies and handle distortions in transition bands. Second, in SBD, a single sub-discriminator accepts the entire 16 sub-bands simultaneously using a convolution layer with input channels 16. On the other hand, in MFBD, a sub-discriminator accepts each sub-band separately using a convolution layer with input channels of 1. However, the parameter of the sub-discriminator is shared across sub-bands as illustrated in Fig. 3(c). Third, in SBD, there exists a sub-discriminator that operates on a transposed sub-band feature. This means that the number of sub-bands becomes the length of the input, and the length of the sub-band feature becomes the channel size of the input. In MFBD, we do not employ such a sub-discriminator. We suspect these overall differences result in the superior performance of MFBD over SBD.

### v2

@@ VI-C Computational Complexity
TABLE IV: Complexity of various neural audio codecs.
@@
We compared HILCodec against other neural audio codecs in terms of computational complexity. We measured the number of parameters, multiply-add counts (MAC) for coding one second long input waveform, and real-time factor (RTF). RTF is defined as the ratio of the length of the input waveform to the time consumed for coding. An RTF higher than 1 means the codec operates in real-time. When calculating RTF, we used one thread of a server CPU (Intel Xeon 6248R 3GHz), exported each model to ONNX Runtime [69], and simulated a streaming condition. Namely, for HILCodec and EnCodec, a chunk with 320 waveform samples was encoded and decoded at a time, and for AudioDec, a chunk with 300 samples was processed at a time. For every convolutional layer with a kernel size larger than one, previous samples were stored in a cache and used to calculate the current chunk. Descript and HiFi-Codec are not streamable, so their RTFs are not reported. The results are shown in Table IV. HILCodec has the smallest number of parameters, requiring minimum memory to store and load model parameters. Moreover, the MAC of HILCodec is smaller than AudioDec, Descript, and HiFi-Codec, and its RTF is higher than 1.
Figure 7: MUSHRA scores with 95% confidence intervals for all audio items.
@@ VII Conclusion
Appendix A Objective Evaluations
Although we have already conducted the MUSHRA listening test, we are sharing the objective evaluations as they may be beneficial to the research community. We measured ViSQOL Audio v3 [67], PEAQ [79], and the 2f-model [80] for different codecs at various bitrates. PEAQ ranges from $-4$ to 00, and the 2f-model ranges from 00 to $100$, with higher scores indicating better quality for both. For PEAQ, we calculated Objective Difference Grade (ODG) using the open-source tool PQevalAudio [^3] http://www-mmsp.ece.mcgill.ca/Documents/Software/ [81]. We calculated the 2f-model scores using a parameter set [^4] https://www.audiolabs-erlangen.de/resources/2019-WASPAA-SEBASS#NewModelParams calibrated for PQevalAudio as given by
$\text{Score}=\frac{56.1345}{1+(-0.0282\times\text{AvgModDiff1}-0.8628)^{2}}\\ -27.1451\times\text{ADB}+86.3515,$ (17)
$\text{FinalScore}=\max(\min(\text{Score},100),0),$ (18)
where AvgModDiff1 and ADB are model output variables from PQevalAudio. As we mentioned in the Section V-D2, we used 300 input audio items at a sampling rate of 24kHz for objective evaluations. All the input and the coded output audios are upsampled to a sampling rate of 48kHz before calculating objective metrics. Since EVS doesn’t support a sampling rate of 24kHz, we upsampled the input audio to 32kHz, compressed with EVS, and upsampled to 48kHz. The results are presented in Table V.
TABLE V: Objective And Subjective Evaluations
Codec (kbps) ViSQOL PEAQ 2f-model MUSHRA
HILCodec (3) 4.128 -3.775 42.897 82.879
Descript (3) 4.149 -3.774 42.431 60.439
EnCodec (3) 4.117 -3.804 41.616 64.364
EVS (5.9) 2.893 -3.805 31.218 61.144
HILCodec (9) 4.407 -3.502 56.282 93.212
Descript (9) 4.414 -3.454 55.743 86.277
EnCodec (9) 4.332 -3.665 50.739 83.742
EVS (16.4) 4.194 -3.663 42.819 86.205
Appendix B MUSHRA Scores For All Items
Since there are only 12 audio items for each of our MUSHRA evaluations, we present MUSHRA scores for all the 12 items in Fig. 7. Note that HILCodec evaluated in (a) and (b) is trained for 468k iterations, while HILCodec evaluated in (c) and (d) is trained for 156k iterations. Some reference items have no confidence interval, indicating unanimous agreement among all assessors on a score of 100.
Appendix C Variance of Residual Branches
@@
In Section III-E2, we construct the VCRB such that the variance after initialization follows (6). Nevertheless, in Fig. 2, although the variance increases linearly with the network depth, it does not strictly follow (6). It is demonstrated that for exactness, we must use scaled weight standardization (Scaled WS) [62]. However, in our early experiments, using Scaled WS resulted in a slightly worse ViSQOL score. Reference [62] also experienced similar degradation and hypothesized that Scaled WS harms the performance of depthwise convolutions. Therefore, as an alternative, we carefully apply initialization for each layer. For layers where an activation function does not come immediately after, we apply LeCun initialization so that $\mathbf{Var}(W\mathbf{x})\approx\mathbf{Var}(\mathbf{x})$ where $W$ is a weight parameter. For layers with subsequent activation functions, we apply He initialization so that $\mathbf{Var}(\sigma(W\mathbf{x}))\approx\mathbf{Var}(\mathbf{x})$. Through these methods, despite the model does not exactly follow (6), it achieves a linear variance increment.
Appendix D Differences Between MFBD and MSD
@@
We present detailed comparisons between SBD used in Avocodo and MFBD used in HILCodec. First, SBD decomposes the input signal into 16 sub-bands while MFBD uses co-prime numbers of sub-bands $N\in\{1,2,3,5,7,11\}$. As described in Section IV-A2, this enables MFBD to model both long- and short-term dependencies and handle distortions in transition bands. Second, in SBD, a single sub-discriminator accepts the entire 16 sub-bands simultaneously using a convolutional layer with input channels 16. On the other hand, in MFBD, a sub-discriminator accepts each sub-band separately using a convolutional layer with input channels of 1. However, the parameter of the sub-discriminator is shared across sub-bands as illustrated in Fig. 3(c). Third, in SBD, there exists a sub-discriminator that operates on a transposed sub-band feature. This means that the number of sub-bands becomes the length of the input, and the length of the sub-band feature becomes the channel size of the input. In MFBD, we do not employ such a sub-discriminator. We suspect these overall differences result in the superior performance of MFBD over SBD.
Appendix E Selection of Audio Samples for MUSHRA
We provide our considerations for selecting audio samples for the MUSHRA experiments. First, we had to choose which audio domains to evaluate. Upon reviewing previous papers, we found that the domains evaluated varied slightly. SoundStream [38] utilized clean speech, noisy speech, noisy and reverberant speech, and music. EnCodec [39] employed clean speech, noisy speech, music, and proprietary music. Descript Audio Codec [42] used clean speech, environmental sounds, and music. We chose to evaluate the most commonly appearing domains: clean speech, noisy speech, and music. Second, we tried to include various sounds. For noisy speech, we included keyboard typing noise, fan noise, the sound of opening chip packets, and the sound of water flowing. The genres of the music samples were pop, soul, electronic, and country rock. Third, for speech codecs, we selected challenging audio samples. We coded audio candidates using an early version of HILCodec, and selected those with large distortions. In our experience, this selection was crucial because the majority of neural speech codecs coded randomly selected speech samples very well, making those samples unsuitable for comparing each model.